# Confronting recent light compact star observations with color-flavor locked quark matter

### v1

@@ 2 Introduction
The recent analysis of Salmi et al. (31) also provided intriguing values for the properties of the pulsar PSR J1231-1411. Notably, their results appeared to be sensitive to the selection of the radius prior they used. By limiting the radius to be consistent with previous observational constraints and nuclear theory, the authors indicated that $M=1.04^{+0.05}_{-0.03}M_{\odot}$ and $R=12.6^{+0.3}_{-0.3}$ km. In addition, by employing an uninformative prior, and constraining the radius between 10 and 14 km, they reported a similar value for $M$ and a higher value for $R$. Interestingly, the recently published work of Qi et al. (32) provided an alternative estimate on the properties of PSR J1231-1411. In particular, by implementing an algorithm based on the spherical star Schwarzschild spacetime and Doppler approximation, the authors inferred that the mass-radius properties of the aforementioned pulsar are $M=1.12^{+0.07}_{-0.07}M_{\odot}$ and $R=9.91^{+0.88}_{-0.86}$ km. In any case, future studies may be required to obtain a more accurate view of this compact object. In the present work, to enrich and extend our investigation, we will consider both estimates of the properties of PSR J1231-1411, referring to the results of Salmi et al. (31) as Case I, and to the results of Qi et al. (32) as Case II.
Given that color-flavor locked (CFL) quark matter (33, 34, 35) has been successfully employed in the reconciliation of HESS J1731-347, not only for its mass and radius but also for its thermal evolution (8, 9), we aim to consider it in order to examine the possible simultaneous explanation of the PSR J1231-1411 properties. To do so, we consider two distinct hypotheses for strange quark matter: a) absolute stability, b) energetically favored at high baryon density. Of utmost importance is to examine if the resulting EOSs satisfy other precisely inferred mass-radius measurements (NICER mission) (36, 37, 38, 39, 40) and also allow for the existence of stable massive stars beyond $2M_{\odot}$ (41, 42, 43). In this way, we aim to provide a general view on the possible existence of CFL quark matter in compact stars by utilizing all available and accurate observational constraints from rotation-powered millisecond pulsars. The main motivation for performing the present work is analyzed in more depth, in the two following paragraphs.
@@
This paper is structured as follows: Section 3 sets the theoretical framework of the present study. Section 4 presents the results for the two distinct hypotheses on the nature of compact stars (quark or hybrid stars) along with a detailed discussion of our findings. Lastly, Section 5 highlights the main insights and conclusions derived from this research.
@@ 3 Color-flavor locked equation of state
The equation of state for CFL quark matter can be formulated within the MIT bag model framework. To the order of $\Delta^{2}$ and $m_{s}^{2}$, where $m_{s}$ represents the mass of the strange quark and $\mu$ the quark chemical potential, the pressure and energy density can be expressed as follows, assuming natural units $(\hbar=c=1)$ (46):
@@
where $a_{4}$ is a parameter that mimics the impact of perturbative QCD (pQCD) corrections (46), $B$ represents the bag constant and $\alpha=-\frac{m_{s}^{2}}{6}+\frac{2\Delta^{2}}{3}$. Notably, an analytical expression for $\varepsilon(P)$ can be obtained by combining the equations presented above:
@@
These formulas reduce to the ones presented in Ref. (45) for $a_{4}=1$, which indicates no strong interactions.
For CFL quark matter to be absolutely stable, its energy per baryon must be lower than the one in the most stable nucleus, ${}^{56}\mathrm{Fe}$ $(m_{b}(Fe)=930\ \mathrm{MeV})$ at zero pressure $(P=0)$ and temperature $(T=0)$ (60). This condition establishes an upper limit for the bag constant as a function of $m_{s}$ and $\Delta$, for a fixed value of $a_{4}$:
@@
This value decreases when considering lower values for $a_{4}$. In this study, we use the constraint derived with $a_{4}=1$, ensuring that all the CFL EOSs used remain absolutely stable even for smaller values of $a_{4}$.
@@
Notably, this relation holds for all of the quark EOSs constructed in the present study.
@@ 4.1 Strange stars
We solved the system of TOV equations (62), using a wide range of $B\ (57-300\ \mathrm{MeV\cdot fm^{-3}}$) and $\Delta\ (50-300\ \mathrm{MeV}$) values to identify the $M-R$ curves that are compatible with the light neutron star measurements for HESS J1731-347 and PSR J1231-1411 (at the $1\sigma$ level). The latter was examined utilizing the results of both Salmi et al. (31) (Case I) and Qi et al. (32) (Case II). Of course, we only considered $B$ and $\Delta$ pairs that lie within the stability window defined by Eq. (4) (fixing the strange quark mass at 95 MeV), while to investigate the possible effects of the $a_{4}$ parameter we employed two distinct values ($a_{4}=1$ and $a_{4}=0.7$). Notably, in the present study we work under the assumption of a universal EOS (a single EOS that accounts for all observations). Therefore, the derived EOSs should be compatible to all state-of-the-art multimessenger constraints on the mass and radius of compact stars. Thus, apart from HESS J1731-347 and PSR J1231-1411, we also utilized the corresponding data related to PSR J0952-0607 (43), PSR J0030+0451 (40) and PSR J0437-4715 (36). Lastly, to highlight the peculiar nature of the recent XTE J1814-338 measurement (63), we have included it in our analysis, although we did not attempt to interpret it simultaneously with the aforementioned constraints.
@@
Figure 1: $\mathrm{B-\Delta}$ pairs that generate $M-R$ curves that are compatible (at the $1\sigma$) with the neutron star observations PSR J0952-0607 (43) (blue region), HESS J1731-347 (7) (yellow region), XTE J1814-338 (63) (red region), PSR J1231-1411 (31) Case I (purple region), PSR J0030+0451 (40) (gray region), PSR J0437-4715 (36) (area between black solid lines) and PSR J1231-1411 (32) Case II (area between purple dashed lines). Panel (a) illustrates the objects’ parameter spaces for Case I and for $a_{4}=1$, while (b) is identical for $a_{4}=0.7$. The right panel, (c), examines Case II for $a_{4}=0.7$. The dark gray area represents the CFL unstable region.
@@
Figure 2: (a) Mass-radius curves that are compatible with PSR J1231-1411 Case I (solid lines) and Case II (dashed lines), with $a_{4}=0.7$. The $B-\Delta$ pairs utilized are representative of the purple region (Case I) and the region between the purple dashed lines (Case II), see Fig. 1(b) and (c). The numbers above each curve represent the $(B,\Delta)$ pairs in $(\mathrm{MeV\cdot fm^{-3},MeV})$. (b) Close-up of PSR J1231-1411’s (31) parameter space considering Case I with $a_{4}=0.7$, plotted together with the maximum masses reached in this window. Notably, a minimum maximum mass of $\sim 3\ \rm{M_{\odot}}$ is needed. (c) Parameter space that is compatible with PSR J0952-0607 (43), PSR J0030+0451 (40), HESS J1731-347 (7), PSR J0437-4715 (36) and PSR J1231-1411 (32) (Case II) measurements, including the maximum masses reached. This area corresponds to the region between the upper limit of PSR J0030+0451 and the lower one of PSR J1231-1411, see Fig. 1(c).
@@
Fig. 1 displays the $(B,\Delta)$ pairs that are compatible with the above measurements for two different values of $a_{4}$, accounting for both estimations of PSR J1231-1411’s properties. The red, purple, blue, yellow and gray parameter spaces correspond to XTE J1814-338, PSR J1231-1411 (Case I), PSR J0952-0607, HESS J1731-347 and PSR J0030+0451 measurements, respectively. The last three areas extend all the way to the horizontal axis, though it is not visible in the graphs. The region between the black solid lines corresponds to PSR J0437-4715, while the one between the purple dashed lines to PSR J1231-1411 (Case II). Fig. 1(a) illustrates the parameter spaces of the above objects for $a_{4}=1$ considering Case I, while Fig. 1(b) shows the same graph for $a_{4}=0.7$. As is evident, although a smaller $a_{4}$ value stiffens the CFL EOS, resulting in a modified stability window, it has only minimal impact on the parameter spaces that are compatible to different observations. Finally, in Fig. 1(c), Case II is examined for $a_{4}=0.7$.
@@
Figs. 1(a) and (b) indicate that HESS J1731-347 and PSR J1231-1411 (Case I), can both be reconciled within the range of parameters that are compatible to the latter, while also respecting the maximum-mass constraints imposed by PSR J0952-0607. This region is characterized by small values for the bag constant, close to the minimum possible one, and large values for the color superconducting gap. As a result, the CFL EOS is extremely stiff, leading to very high maximum masses for all $(B,\Delta)$ pairs (above $3\ M_{\odot}$). The latter is demonstrated in Fig. 2(b), where the maximum masses for all $(B,\Delta)$ combination of the purple region of Fig. 1(b) are calculated (for the $a_{4}=0.7$ case).
@@
Another interesting point from Figs. 1(a) and (b) is related to the fact that the parameter spaces which are compatible with PSR J0437-4715 and PSR J1231-1411 (Case I) do not overlap at the $1\sigma$ level, implying that they cannot be simultaneously explained within this model. However, this tension is very mild, as we have checked that a narrow common parameter space exists, when considering the $2\sigma$ estimation for PSR J0437-4715 (see Fig. 2(a), where we have plotted the $M-R$ diagrams for representative EOSs which are compatible with PSR J1231-1411).
@@
Considering Case II, Fig. 1(c) shows that there is a small range of parameters which is compatible with all measurements at the $1\sigma$ level (of course except XTE J1814-338). This area is defined by the lower boundary from PSR J1231-1411 (purple dashed line) and the upper limit from PSR J0030+0451 (gray region). Notably, for PSR J0030+0451 we employed the measurement of Miller et al. (40). The aforementioned common areas do not change significantly when adopting the measurements of Riley et al. (39), and it would increase when using the lower (regarding the radius) of the three estimations provided by Vinciguerra et al. (38).
@@
The parameter space in which all measurements can be explained by pure CFL matter (Case II) is shown in Fig. 2(c), alongside the corresponding maximum mass predictions. The resulting maximum masses span the range of approximately $2.25-3.25\ \rm{M_{\odot}}$, which is consistent with current observations of heavy pulsars. This maximum mass range, along with that obtained for the parameter space compatible with PSR J1231-1411 (Case I) for $a_{4}=0.7$, points toward the possible existence of quark stars with masses exceeding $3\ \rm{M_{\odot}}$. Notably, such masses reside within the so-called lower mass gap (64, 65), in which the nature of compact objects remains undetermined. Interestingly, evidence for compact objects within the mass gap has emerged from multiple recent studies, based on observations of non-interacting binary systems (66, 67), radio pulsar surveys (68), and gravitational wave detections from binary mergers (69, 70, 71, 72, 73).
@@
A final remark needs to be made with regards to the possible explanation of the thermal evolution for the CCO in the HESS J1731-347 remnant. Notably, the CCO has a rather high temperature for its estimated age which suggests slow cooling (8), analogous to those of purely hadronic stars. Previous works (8, 9) have qualitatively suggested that superconductivity may suppress the rapid cooling processes that are expected in unpaired strange quark matter and therefore the temperature of the CCO could be explained. Interestingly, according to Ref. (9) this could occur only under the consideration of vanishingly small gap values $\Delta<1$ MeV. This is related to the fact that for large values of pairing gaps the suppression of standard cooling processes in the stellar core is dramatic. Therefore, there is a potential tension between the necessity for large gaps, dictated by the large maximum mass constraints, and the low gaps suggested by qualitative analyses on the cooling of the CCO in HESS J1731-347. In addition, it remains an open question whether the low central density expected in the low mass CCO is sufficient for CFL matter to appear for vanishingly small pairing gaps (considering the conditions that need to be met for CFL matter to be energetically favorable (44, 61)). In any case, a future precise analysis on the cooling of CFL stars incorporating all of the aforementioned issues, including also potential uncertainties on the estimated temperature and age, is essential to safely conclude about whether the CCO in the HESS J1731-347 could be in fact a CFL strange star.

### v2

@@ 2 Introduction
The recent analysis of Salmi et al. (31) also provided intriguing values for the properties of the pulsar PSR J1231-1411. Notably, their results appeared to be sensitive to the selection of the radius prior they used. By limiting the radius to be consistent with previous observational constraints and nuclear theory, the authors indicated that $M=1.04^{+0.05}_{-0.03}M_{\odot}$ and $R=12.6^{+0.3}_{-0.3}$ km. In addition, by employing an uninformative prior, and constraining the radius between 10 and 14 km, they reported a similar value for $M$ and a higher value for $R$. Interestingly, the recently published work of Qi et al. (32) provided an alternative estimate on the properties of PSR J1231-1411. In particular, by implementing an algorithm based on the spherical star Schwarzschild spacetime and the Doppler approximation, the authors inferred that the mass-radius properties of the aforementioned pulsar are $M=1.12^{+0.07}_{-0.07}M_{\odot}$ and $R=9.91^{+0.88}_{-0.86}$ km. In any case, future studies may be required to obtain a more accurate view of this compact object. In the present work, to enrich and extend our investigation, we will consider both estimates of the properties of PSR J1231-1411, referring to the results of Salmi et al. (31) as Case I, and to the results of Qi et al. (32) as Case II.
Given that color-flavor locked (CFL) quark matter (33, 34, 35) has been successfully employed in the reconciliation of HESS J1731-347, not only for its mass and radius but also for its thermal evolution (8, 9), we aim to consider it in order to examine the possible simultaneous explanation of the PSR J1231-1411 properties. To do so, we consider two distinct hypotheses for strange quark matter (SQM): a) absolutely stable, b) energetically favored at high baryon density. Of utmost importance is to examine if the resulting EOSs satisfy other precisely inferred mass-radius measurements (NICER mission) (36, 37, 38, 39, 40) and also allow for the existence of stable massive stars beyond $2M_{\odot}$ (41, 42, 43). In this way, we aim to provide a general view on the possible existence of CFL quark matter in compact stars by utilizing all available and accurate observational constraints from rotation-powered millisecond pulsars. The main motivation for performing the present work is analyzed in more depth in the two following paragraphs.
@@
This paper is structured as follows: Section 3 sets the theoretical framework of the present study. Section 4 presents the results for the two distinct hypotheses on the nature of compact stars (quark or hybrid stars) along with a detailed discussion on our findings. Lastly, Section 5 highlights the main insights and conclusions derived from this research.
@@ 3 Color-flavor locked equation of state
The equation of state for CFL quark matter can be formulated within the MIT bag model framework. To the order of $\Delta^{2}$ and $m_{s}^{2}$, where $m_{s}$ represents the mass of the strange quark and $\mu$ the quark chemical potential, the pressure and energy density can be expressed as follows $(\hbar=c=1)$ (45, 46):
@@
where $a_{4}$ is a parameter that mimics the impact of perturbative QCD (pQCD) corrections (46), $B$ represents the bag constant and $\alpha=-\frac{m_{s}^{2}}{6}+\frac{2\Delta^{2}}{3}$. Notably, an analytical expression for $\varepsilon(P)$ can be obtained by combining Eqs. (1) and (2):
@@
Eq. (3) reduces to the one presented in Ref. (45) for $a_{4}=1$, which indicates no strong interactions.
For CFL quark matter to be absolutely stable, its energy per baryon must be lower than the one in the most stable nucleus, ${}^{56}\mathrm{Fe}$ $(m_{b}(Fe)=930\ \mathrm{MeV})$ at zero pressure and temperature (60). This condition establishes an upper limit for the bag constant as a function of $m_{s}$ and $\Delta$, for a fixed value of $a_{4}$:
@@
This value decreases when considering lower values for $a_{4}$. In this study, we use the constraint derived with $a_{4}=1$ (Eq. (5)), ensuring that all the CFL EOSs used characterize absolutely stable quark matter.
@@
Notably, Eq. (6) holds for all of the quark EOSs constructed in the present study.
@@ 4.1 Strange stars
In this section, we present our results derived under the assumption that SQM represents the true ground state of matter. However, a key question, that requires a proper discussion, immediately arises: if SQM is more stable than nuclear matter why does normal matter persist? One explanation may be that nuclear matter is metastable (see Refs. (62, 63, 64, 65, 66, 67) and references therein), separated from the favorable SQM state by a significant energy barrier. Thus, at low densities, the conversion to SQM may be suppressed due to quantum tunneling limitations. However, the extreme neutron star environment, characterized by high densities, may enhance the possibility of a SQM droplet appearing via quantum fluctuations. In addition, extreme events such as supernovae and neutron star mergers may also facilitate SQM formation. Given that the appearance of a SQM seed could trigger the conversion of the entire hadronic star into a strange star, it has long been hypothesized that all compact stars may be of SQM nature.
@@
Figure 1: $\mathrm{B-\Delta}$ pairs that generate $M-R$ curves that are compatible (at the $1\sigma$ level) with the neutron star observations PSR J0952-0607 (43) (blue region), HESS J1731-347 (7) (yellow region), XTE J1814-338 (70) (red region), PSR J1231-1411 (31) Case I (purple region), PSR J0030+0451 (40) (gray region), PSR J0437-4715 (36) (area between black solid lines) and PSR J1231-1411 (32) Case II (area between purple dashed lines). (a) illustrates the parameter spaces when considering Case I and for $a_{4}=1$, while (b) is identical for $a_{4}=0.7$. (c,d) examine Case II for $a_{4}=1$ and $a_{4}=0.7$ respectively. The dark gray area represents the region where CFL quark matter is not absolutely stable.
Proceeding with our calculations, we solved the system of the TOV equations (68), using a wide range of $B\ (57-300\ \mathrm{MeV\cdot fm^{-3}}$) and $\Delta\ (50-300\ \mathrm{MeV}$) values to identify the $M-R$ curves that are compatible with the light neutron star measurements for HESS J1731-347 and PSR J1231-1411 (at the $1\sigma$ level). The latter was examined utilizing the results of both Salmi et al. (31) (Case I) and Qi et al. (32) (Case II). Of course, we only considered $B$ and $\Delta$ pairs that lie within the stability window defined by Eq. (4) (fixing the strange quark mass at 95 MeV), while to investigate the possible effects of the $a_{4}$ parameter we employed two distinct values, $a_{4}=1$ and $a_{4}=0.7$. The latter value was selected as it results to EOSs similar to the quark matter EOS from pQCD (e.g., see Refs. (46, 69)).
Notably, in the present study we work under the assumption of a universal EOS (a single EOS that accounts for all observations). Therefore, the derived EOSs should be compatible to all state-of-the-art multimessenger constraints on the mass and radius of compact stars. Thus, apart from HESS J1731-347 and PSR J1231-1411, we also utilized the corresponding data related to PSR J0952-0607 (43), PSR J0030+0451 (40) and PSR J0437-4715 (36). Lastly, to highlight the peculiar nature of the recent XTE J1814-338 measurement (70), we have included it in our analysis, although we did not attempt to interpret it simultaneously with the aforementioned constraints (for some recent works on its reconciliation see Refs. (71, 72, 73, 74, 75)).
@@
Figure 2: (a) Mass-radius curves that are compatible with PSR J1231-1411 Case I (solid lines) and Case II (dashed lines), with $a_{4}=0.7$. The $B-\Delta$ pairs utilized are representative of the purple region (Case I) and the region between the purple dashed lines (Case II), see Fig. 1(b) and (d). The numbers above each curve represent the $(B,\Delta)$ pairs in $(\mathrm{MeV\cdot fm^{-3},MeV})$. (b) Close-up of PSR J1231-1411’s (31) parameter space considering Case I with $a_{4}=0.7$, plotted together with the maximum masses reached in this window. Notably, a minimum maximum mass of $\sim 3\ \rm{M_{\odot}}$ is needed. (c) Parameter space that is compatible with PSR J0952-0607 (43), PSR J0030+0451 (40), HESS J1731-347 (7), PSR J0437-4715 (36) and PSR J1231-1411 (32) (Case II) measurements, including the maximum masses reached. This area corresponds to the region between the upper limit of PSR J0030+0451 and the lower one of PSR J1231-1411, see Fig. 1(d).
@@
Fig. 1 displays the $(B,\Delta)$ pairs that are compatible with the above measurements for two different values of $a_{4}$, accounting for both estimations of PSR J1231-1411’s properties. The red, purple, blue, yellow and gray parameter spaces correspond to XTE J1814-338, PSR J1231-1411 (Case I), PSR J0952-0607, HESS J1731-347 and PSR J0030+0451 measurements, respectively. The last three areas extend all the way to the horizontal axis, though it is not visible in the graphs. The region between the black solid lines corresponds to PSR J0437-4715, while the one between the purple dashed lines to PSR J1231-1411 (Case II). Fig. 1(a) illustrates the parameter spaces of the above objects for $a_{4}=1$ considering Case I, while Fig. 1(b) shows the same graph for $a_{4}=0.7$. As is evident, although a smaller $a_{4}$ value alters the CFL EOS, resulting in a modified stability window, it has only minimal impact on the parameter spaces that are compatible to different observations. Finally, in Figs. 1(c) and (d), Case II is examined for $a_{4}=1$ and $a_{4}=0.7$ respectively.
@@
Figs. 1(a) and (b) indicate that HESS J1731-347 and PSR J1231-1411 (Case I), can both be reconciled within the range of parameters that are compatible to the latter, while also respecting the maximum-mass constraints imposed by PSR J0952-0607. This region is characterized by small values for the bag constant, close to the minimum possible one, and large values for the color superconducting gap. As a result, the CFL EOS is extremely stiff, leading to very high maximum masses for all $(B,\Delta)$ pairs (above $3\ M_{\odot}$). The latter is demonstrated in Fig. 2(b), where the maximum masses for all $(B,\Delta)$ combinations of the purple region of Fig. 1(b) are calculated (for the $a_{4}=0.7$ case).
@@
Another interesting point from Figs. 1(a) and (b) is related to the fact that the parameter spaces which are compatible with PSR J0437-4715 and PSR J1231-1411 (Case I) do not overlap at the $1\sigma$ level, implying that they cannot be simultaneously explained within this model. However, this tension is mild, as we have checked that a narrow common parameter space exists, when considering the $2\sigma$ estimation for PSR J0437-4715 (see Fig. 2(a), where we have plotted the $M-R$ diagrams for representative EOSs which are compatible with PSR J1231-1411).
@@
Considering Case II, Figs. 1(c) and (d) show that there is a small range of parameters which is compatible with all measurements at the $1\sigma$ level (of course except XTE J1814-338). These areas are defined by the lower boundary from PSR J1231-1411 (purple dashed line) and the upper limit from PSR J0030+0451 (gray region). Notably, for PSR J0030+0451 we employed the measurement of Miller et al. (40). The aforementioned common areas do not change significantly when adopting the measurements of Riley et al. (39), and they would increase when using the lower (regarding the radius) of the three estimations provided by Vinciguerra et al. (38).
@@
The parameter space in which all measurements can be explained by pure CFL matter (Case II) is shown in Fig. 2(c), alongside the corresponding maximum mass predictions (for the $a_{4}=0.7$ case). The resulting maximum masses span the range of approximately $2.25-3.25\ \rm{M_{\odot}}$, which is consistent with current observations of heavy pulsars. This maximum mass range, along with that obtained for the parameter space compatible with PSR J1231-1411 (Case I) for $a_{4}=0.7$, points toward the possible existence of quark stars with masses exceeding $3\ \rm{M_{\odot}}$. Notably, such masses reside within the so-called lower mass gap (76, 77), in which the nature of compact objects remains undetermined. Interestingly, evidence for compact objects within the mass gap has emerged from multiple recent studies, based on observations of non-interacting binary systems (78, 79), radio pulsar surveys (80), and gravitational wave detections from binary mergers (81, 82, 83, 84, 85).
@@
A final remark needs to be made with regards to the possible explanation of the thermal evolution for the CCO in the HESS J1731-347 remnant. Notably, the CCO has a rather high temperature for its estimated age which suggests slow cooling (8), analogous to those of purely hadronic stars. Previous works (8, 9) have qualitatively suggested that superconductivity may suppress the rapid cooling processes that are expected in unpaired SQM and therefore the temperature of the CCO could be explained. Interestingly, according to Ref. (9) this could occur only under the consideration of vanishingly small gap values $\Delta<1$ MeV. This is related to the fact that for large values of pairing gaps the suppression of standard cooling processes in the stellar core is dramatic. Therefore, there is a potential tension between the necessity for large gaps, dictated by the large maximum mass constraints, and the low gaps suggested by qualitative analyses on the cooling of the CCO in HESS J1731-347. In addition, it remains an open question whether the low central density expected in the low mass CCO is sufficient for CFL matter to appear for vanishingly small pairing gaps (considering the conditions that need to be met for CFL matter to be energetically favorable (44, 61)). In any case, a future precise analysis on the cooling of CFL stars incorporating all of the aforementioned issues, including also potential uncertainties on the estimated temperature and age, is essential to safely conclude about whether the CCO in the HESS J1731-347 could be in fact a CFL strange star.